OpenStudy (anonymous):

Tickets to a concert cost $2 for children and $5 for adults. If 630 people attended the concert and the ticket sales totaled $2,910, how many children attended the concert?

OpenStudy (jdoe0001):

a = amount of adults c = amount of children 5 bucks per adult, sales for that will be 5 * a, or 5a 2 bucks per child, sales for that will be 2 * c, or 2c so 630 people attended, adults and children summed up will be \(\bf \color{blue}{a + c = 630}\) so total sales was $2,910 so adult sales and children sales summed up will be \(\bf \color{blue}{5a + 2c = 2910}\) ------------------------------------------------ from a + c = 630 we could solve for "a" and say that \(\bf a=\color{red}{ 630 - c}\) so now we know that a = 630-c, so let us use that in the 2nd equation \(\bf \color{blue}{5(\color{red}{630-c}) + 2c = 2910}\) solve for "c" to find out how many children attended

OpenStudy (anonymous):

Thank you @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

It was 80, so thank you again. Though my test crashed due to poor internet.